OpenStudy (anonymous):

the sum of a number and its reciprocal is 13. whats this number

jimthompson5910 (jim_thompson5910):

You have to solve this equation \[\Large x + \frac{1}{x} = 13\]

jimthompson5910 (jim_thompson5910):

I would multiply everything by the LCD x to get \[\Large x + \frac{1}{x} = 13\] \[\Large x^2 + 1 = 13x\] then get everything to one side \[\Large x^2 + 1 = 13x\] \[\Large x^2-13x + 1 = 0\] from here, use the quadratic formula to solve for x

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

you're welcome